# OM4OV: Leveraging Ontology Matching for Ontology Versioning

### v1

@@ 3.4. Pipeline Optimisation
Figure 3 illustrates the cross-reference mechanism used in the OM4OV pipeline. We can see that, without using the cross-reference $O_{r}$, the matching candidates cover the range of $O\cup O^{\prime}$. This number can be significantly reduced by removing prior matches (i.e. $O\cap O_{r}\cap O^{\prime}$) and non-matches (i.e. $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$). The prior matching will be part of the final alignment, while the known non-matching will be completely removed in the subsequent OV process. The OV process then only determines the posterior alignment. In practice, the prior alignment usually contains a large number of remain entities and a small number of update entities. Matching performance is also improved by using these known mappings. Since $A_{\pi}$ are inferred from the OM references validated by domain experts, they represent a solid ground truth for alignment in a specific domain. On the other hand, while the known non-matches are removed from the OV process, this reduces the complexity of detecting the posterior alignment.
Figure 3. Cross-reference mechanism.
@@
    (11) $A_{\pi}=R_{or}\cap R_{o^{\prime}r}=\{(e1,e2,\equiv,c)|e1,e2\in O\cap O_{r}\cap O^{\prime},c\geq s\}$
  2. (2)
@@
    (12) $A_{\pi^{*}}=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{r},e2\in O^{\prime}\setminus O_{r},c\geq s\}$
@@
(13) $A_{\pi}=(R_{or1}\cap R_{o^{\prime}r1})\cup(R_{or2}\cap R_{o^{\prime}r2})\cup...\cup(R_{orn}\cap R_{o^{\prime}rn})$
@@
Our cross-reference mechanism has no impact on our proposed measures for OV. However, our cross-reference mechanism is incorporated into our OV testbed. For this, $R_{or}$ and $R_{o^{\prime}r}$ are created according to the following rules:
  1. (1)
    $R_{or}$ is the original reference.xml removing all the mappings related to add entities.
  2. (2)
    $R_{o^{\prime}r}$ is the original reference.xml removing all the mappings related to delete entities and updating all the mappings related to update entities.
4. Implementation & Evaluation
The OM4OV pipeline is implemented in Agent-OV, a variant of Agent-OM (Qiang et al., 2024). Agent-OM is an agent-powered LLM-based OM system. Its foundation framework is designed for traditional OM tasks. We extend the original framework with the OM4OV pipeline so that it can be used to handle OV tasks.
We use the LLM model gpt-4o-mini for evaluation. As an extension of Agent-OM, Agent-OV also supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT (OpenAI, 2024), Anthropic Claude (Anthropic, 2024), and Mistral AI (Mistral AI, 2024), as well as open-source LLMs Meta Llama (Meta, 2024), Google Gemma (Gemma Team, Google DeepMind, 2024), and Alibaba Qwen (Alibaba, 2024). For the performance of Agent-OV using different LLMs, we refer the reader to (Qiang et al., 2024), where we find that API-accessed LLMs generally perform better than open-source LLMs. The hyperparameter settings are set to $similarity\_threshold=0.90$ and $top@k=3$ across all alignments generated from the OV testbed.
4.1. Evaluation of OV testbed
The current version of the OV testbed contains a total of 12 distinct ontologies from three different OAEI tracks. Table 1 lists the detailed information of the selected OAEI track used for the OV testbed. The anatomy track contains two large ontologies, while the MSE track has three medium ontologies, and the conference track has 7 small ontologies.
Table 1. Selected OAEI track for the OV testbed.
Figures 45, and 6 show the evaluation of the Agent-OV system on the OV testbed. The results indicate that it is possible to unify the OM and OV tasks. With the necessary modifications, the OM systems can also be used in OV tasks.
  1. (1)
@@ 4.1. Evaluation of OV testbed
      Figure 4. Evaluation of Anatomy Track.
      Figure 5. Evaluation of Conference Track.
      Figure 6. Evaluation of MSE Track.
@@
4.2. Evaluation of pipeline optimisation
We apply the cross-reference mechanism on the same alignments that we evaluated in Section 4.1. Figure 7 shows the comparison of the OV performance with and without using the cross-reference (CR) mechanism to detect update entities. We can see that the CR mechanism significantly improves recall and precision in most ontologies, resulting in a solid improvement in the F1 score.
(a) Precision
(b) Recall
(c) F1 Score
Figure 7. Evaluation of cross-reference (CR) mechanism.
4.3. Evaluation of hyperparameter settings
Unlike OM tasks, changes in hyperparameter settings in OV tasks do not lead to a trade-off in precision and recall. Instead, the hyperparameter settings directly influence the sub-measures. Lower similarity thresholds and higher top@k values can result in more update entities being detected, while higher similarity thresholds and lower top@k values may find more add and delete entities. The hyperparameter setting has no impact on remain entities.
5. Discussion
So far, we have experimentally validated the OM4OV pipeline with a novel cross-reference mechanism for pipeline optimisation. Although matching performance has been improved, it still does not achieve 100% as false mappings may still exist. Are these mappings genuinely false?
  1. (1)
    False mappings in OV can be a wrong ontology design choice. In the following example, the reference shows that cmt:writtenBy is updated to cmt:isWrittenBy, but the OV system will consider this entity to remain unchanged because a false mapping (cmt:writtenBy, cmt:writtenBy, $\equiv$, 1) is detected in the remain subset. This issue is caused by incorrect name conventions. In the CMT ontology, cmt:hasAuthor and cmt:writtenBy are different entities because one is targeting the paper (cmt:Paper, cmt:hasAuthor, cmt:Author) and one is targeting the review (cmt:review, cmt:writtenBy, cmt:reviewer). However, using cmt:writtenBy for paper still makes sense (cmt:Paper, cmt:writtenBy, cmt:Author). Within one ontology, the meaning of two entities is too close to be distinguished and therefore leads to them being synonyms for each other. Ideally, we should avoid this type of ontology design. This example also demonstrates a unique benefit of using OM for OV, which could potentially aid in a more effective ontology design.
@@ 5. Discussion
  3. (3)
    False mappings in OV can be an inappropriate setup on the similarity threshold. In the following example, if $similarity\_threshold=0.95$, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If $similarity\_threshold=0.90$, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$. Both results are valid because defining the boundary between matching and non-matching is context- and application-dependent. For example, the similarity threshold is relatively higher in the biomedical domain to ensure each terminology is unique, whereas the similarity threshold in the conference domain can be relatively lower to improve the interoperability of terminologies used in research conferences.
    similarity_threshold = 0.95, cmt:SubjectArea => None
    similarity_threshold = 0.90, cmt:SubjectArea => cmt:Topic
6. Limitations
  1. (1)
    We focus mainly on tracking conceptual changes of classes and properties (e.g. adding, deleting, or updating a class). In practice, there are also internal relationship changes (e.g. changing the domain and range of a class or moving a sibling class to a different parent). These changes are currently only indicated by the changes in the similarity score, while details can only be observed by inspecting the classes and properties. Our future work aims to improve the explainability of changes.
  2. (2)
@@ 7. Conclusion
Our approach is compatible with ontologies using different versioning methods (using URIs or additional versioning triples), and even ontologies missing or without version information. Our approach stores the version information independently from the ontology, offering a simple and lightweight way to track versioning changes in ontologies.
Acknowledgements.
The authors thank Weiqing Wang from Monash University for giving helpful advice on this work. The authors also thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.

### v2

@@ III-D Cross-Reference Mechanism
Fig. 3 illustrates the cross-reference mechanism used in the OM4OV pipeline. We can see that, without using the cross-reference $O_{r}$, the matching candidates cover the range of $O\cup O^{\prime}$. This number can be significantly reduced by removing prior matches (i.e. $O\cap O_{r}\cap O^{\prime}$) and non-matches (i.e. $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$). The prior matching will be part of the final alignment, while the known non-matching will be completely removed in the subsequent OV process. The OV process then only determines the posterior alignment. In practice, the prior alignment usually contains a large number of remain entities and a small number of update entities. Matching performance is also improved by using these known mappings. Since prior matches are inferred from the OM references validated by domain experts, they represent a solid ground truth for alignment in a specific domain. On the other hand, while the known non-matches are removed from the OV process, this reduces the complexity of detecting the posterior alignment.
Figure 3: Cross-reference mechanism.
@@
    $A_{\pi}=R_{or}\cap R_{o^{\prime}r}=\{(e1,e2,\equiv,c)|e1,e2\in O\cap O_{r}\cap O^{\prime},c\geq s\}$ (12)
  2. 2.
@@
    $A_{\pi^{*}}=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{r},e2\in O^{\prime}\setminus O_{r},c\geq s\}$ (13)
@@
$A_{\pi}=(R_{or1}\cap R_{o^{\prime}r1})\cup(R_{or2}\cap R_{o^{\prime}r2})\cup...\cup(R_{orn}\cap R_{o^{\prime}rn})$ (14)
@@
IV Implementation & Evaluation
The OM4OV pipeline is implemented in Agent-OV, a variant of Agent-OM [14]. Agent-OM is an agent-powered LLM-based OM system. Its foundation framework is designed for traditional OM tasks. We extend the original framework with the OM4OV pipeline so that it can be used to handle OV tasks.
As an extension of Agent-OM, Agent-OV also supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT [15], Anthropic Claude [16], and Mistral AI [17], as well as open-source LLMs Meta Llama [18], Google Gemma [19], and Alibaba Qwen [20]. For the performance of Agent-OV using different LLMs, we refer the reader to [14], where we find that API-accessed LLMs generally perform better than open-source LLMs.
IV-A Evaluation of OV testbed
Fig. 4 shows the evaluation of the Agent-OV system on the OV testbed. We use the OpenAI model gpt-4o-mini (2024-07-18) [21] for evaluation. The embedding chosen correspondingly is OpenAI text-embedding-ada-002 [22]. The hyperparameter settings are $similarity\_threshold=0.90$ and $top@k=3$ across all alignments generated. We run the OV testbed for 5 iterations, simulating 5 distinct versions for each ontology. The results indicate that it is possible to unify the OM and OV tasks. With the necessary modifications, the OM systems can also be used in OV tasks.
(a) Anatomy Track: human
(b) Anatomy Track: mouse
(c) Conference Track: cmt
(d) Conference Track: conference
(e) Conference Track: confof
(f) Conference Track: edas
(g) Conference Track: ekaw
(h) Conference Track: iasted
(i) Conference Track: sigkdd
(j) MSE Track: EMMO
(k) MSE Track: MatInfo
(l) MSE Track: MatOnto
Figure 4: Evaluation of OV testbed.
  1. 1.
@@ IV-A Evaluation of OV testbed
IV-B Evaluation of cross-reference mechanism
Fig. 5 shows the comparison of the OV performance with and without using the cross-reference (CR) mechanism on the same OV testbed we analysed in Section IV-A. We can see that the CR mechanism significantly improves the capability of Agent-OV to detect update entities and slightly improves the performance on detecting add and delete entities, with no effect on the remain entities. More specifically, the performance improvement is mainly caused by the recall for update entities, while this is more evident in the precision for add and delete entities.
(a) Remain-Precision
(b) Update-Precision
(c) Add-Precision
(d) Delete-Precision
(e) Remain-Recall
(f) Update-Recall
(g) Add-Recall
(h) Delete-Recall
(i) Remain-F1 Score
(j) Update-F1 Score
(k) Add-F1 Score
(l) Delete-F1 Score
Figure 5: Evaluation of cross-reference (CR) mechanism.
IV-C Evaluation of hyperparameter settings
Fig. 6 shows the comparison of using different similarity thresholds in the CMT Ontology versioning. Similar to OM tasks, changes in hyperparameter settings in OV tasks will lead to a trade-off in precision and recall. Moreover, the hyperparameter settings also influence the sub-measures. We find that the similarity threshold has no effect on detecting remain entities, but significantly affects the detection of update entities and slightly affects the detection of add and delete entities. Also, detecting update entities and detecting add and delete is negatively correlated. Lower similarity thresholds can result in more update entities being detected, while higher similarity thresholds may find more add and delete entities. The result of the experiment also shows that our cross-reference mechanism is less sensitive to the similarity threshold, which means that it can be useful in scenarios where the optimal similarity threshold is unclear or hard to determine.
Figure 6: Evaluation of different similarity thresholds.
V Discussion
So far, we have experimentally validated the OM4OV pipeline with a novel cross-reference mechanism for pipeline optimisation. Although matching performance has been improved, it still does not achieve 100% as false mappings may still exist. Are these mappings actually false?
  1. 1.
    False mappings in OV can be a wrong ontology design choice. In the following example, the reference shows that cmt:writtenBy is updated to cmt:isWrittenBy, but the OV system will consider this entity to remain unchanged because a false mapping (cmt:writtenBy, cmt:writtenBy, $\equiv$, 1) is detected in the remain subset. This issue is caused by incorrect name conventions. In the CMT ontology, cmt:hasAuthor and cmt:writtenBy are different entities because one is targeting the paper (cmt:Paper, cmt:hasAuthor, cmt:Author) and one is targeting the review (cmt:review, cmt:writtenBy, cmt:reviewer). However, using cmt:writtenBy for paper still makes sense (cmt:Paper, cmt:writtenBy, cmt:Author). Within one ontology, the meaning of two entities is too close to be distinguished and therefore leads to them being synonyms for each other. Ideally, we should avoid this type of ontology design. This example also demonstrates a unique benefit of using OM for OV, which could potentially assist in ontology design.
@@ V Discussion
  3. 3.
    False mappings in OV can be an inappropriate setup on the similarity threshold. In the following example, if $similarity\_threshold\ (s)=0.95$, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If $similarity\_threshold\ (s)=0.90$, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$. Both results are valid because defining the boundary between matching and non-matching is context- and application-dependent. For example, the similarity threshold is relatively higher in the biomedical domain to ensure each terminology is unique, whereas the similarity threshold in the conference domain can be relatively lower to improve the interoperability of terminologies used in research conferences.
    s = 0.95, cmt:SubjectArea => None
    s = 0.90, cmt:SubjectArea => cmt:Topic
VI Limitations
We focus mainly on tracking conceptual changes in classes and properties (e.g. adding, deleting, or updating). In practice, there are also internal relationship changes (e.g. changing the domain and range of a class or moving a sibling class to a different parent). These changes are currently only indicated by the changes in the similarity score, while details can only be observed by inspecting the classes and properties. Our future work aims to improve the explainability of changes.
@@ VII Conclusion
Our approach is compatible with ontologies using different versioning methods (using URIs or additional versioning triples), and even ontologies missing or without version information. Our approach stores the version information independently from the ontology, offering a simple and lightweight way to track versioning changes in ontologies. In the future, we plan to apply the OM4OV pipeline in emerging domain ontologies that require consistent changes over time (e.g. smart building ontologies Brick Schema [23], RealEstateCore [24], and ASHRAE Standard 223P [25]).
Supplemental Material Statement
Agent-OV is an extended version of Agent-OM. Agent-OM can be downloaded from https://github.com/qzc438-research/ontology-llm. The OAEI track repositories used in this study can be downloaded from OAEI MELT at https://dwslab.github.io/melt/track-repository (date accessed: 2024-06-01). The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
Acknowledgment
The authors would like to thank Alice Richardson of the Statistical Support Network, Australian National University (ANU), for helpful advice on the statistical analysis in this paper. The authors also thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.